Exercise 3. Let u(t) denote the unit step function u(t)= { 0, t<0, u(t)={ 1, ost. Show that A) (exp(at) u(t)) * (exp(Bt) u(t)Please do all parts please. thank you
Answer #1

• In a linear hamon ... antenna 1 act) unit step function, UH) =r otko 11i ost @de t store 104) * 234) — — Xi(s).x265) at ayat Bt e ult* é OCH at Bt е – е 2-BC A+B=0 A--B7-) = array satult) ulet) = A termes 18-a) (S-B) = (4-p) - (-p) (5-2) (s-B) AppSa (9-~)*s-s) (S-) . 1 EACS-«)(S-1)+ BCS--B) + C CS-x) 1= A[s-sp-s« +] + BS-BB + c(sta +254] 1- M+c]s+ -AB-A+B-266]stad Bsential and Apply Inverse Laplace transform tept**Jur). (X-B) © testuct) * et cuch) = (8- ) foomula of 67. 65-633 - Luces = aul) © t >* i Stav) = (cowym sose cal (S-2) .. 31-34) fame te eeu) 13 tiene storia éta Inte) 1 tsett utt). - ans c estu(-4) *cerpes Casa EFB - S-2H-B) = A(S-B) + BCS-2) 18-) 456) -1 = AS - AB +B5-B2. ~ = (A+B)s - AB-Bé A+B=0 -AB-BX = -1 AB+BP = 1 ß (get ul-t) * ee fuckt); if Re(@)Repy seda eties . ang e Real If Recas Relb), there is NO comomon Region of convergenc croc),